OpenStudy (anonymous):

7th grade math plz @paki

OpenStudy (anonymous):

Estimate the value of this expression by rounding each number to the nearest whole unit. (1.94)(7.51)(3.74) A. 56 B. 64 C. 21 D. 42

OpenStudy (anonymous):

Rounding each to the nearest whole number makes it (2x8x4)

OpenStudy (anonymous):

so can u explain more please like walk me t hrough it plz

OpenStudy (k_lynn):

what would you round all of those numbers to?

OpenStudy (k_lynn):

like 1.94 would be closest to what. 7.51 is closest to what. and 3.74 is closest to what whole number?

OpenStudy (anonymous):

1.94 is closer to 2; 7.51 is close to 8; 3.74 is closer to 4; so 2 * 8 * 4

OpenStudy (anonymous):

2 * 8 = 16 * 4 = 64

OpenStudy (anonymous):

so ur answer is B